# Failure Modes of Maximum Entropy RLHF

### v1

@@ Appendix B Margins and Overoptimization
It has been shown that methods such as SimPO can achieve performance comparable to DPO even with a target margin of $\gamma=0$, as demonstrated in the original SimPO paper. This suggests that offline methods do not necessarily require reference models when operating within the safe KL region, and that introducing margins generally improves performance across benchmarks. This effect arises from both model capabilities and dataset coverage: larger models are less prone to common overfitting behaviors and can extract more meaningful signals during optimization, rather than engaging in reward hacking, a phenomenon observed in both online and offline preference optimization (Gao et al., 2022; Rafailov et al., 2024c). Consequently, the influence of the reference model is minimal and can often be neglected. However, this behavior is contingent on the task being sufficiently challenging and the model being strong enough to avoid overoptimization. To validate this observation, we train Pythia-1B on TL;DR using SimPO across different margin values ($\gamma$) and learning rates, in a setting where the model is relatively weaker and the task is easier compared to standard chat datasets such as UltraFeedback (Cui et al., 2024) used in SimPO.
@@
Appendix C Extra Figures
Appendix D Mathematical Derivations for Maximum Entropy RL
D.1 Deriving the Optimum of the Entropy-Regularized Reward Maximization Objective
In this appendix, we will derive the optimal policy for Maximum Entropy RL. Analogously to the KL-constrained case (Rafailov et al., 2024c), we optimize the following objective:
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi}\bigl[r(x,y)\bigr]+\alpha\mathcal{H}\bigl[\pi(y|x)\bigr]$ (19)
@@ D.1 Deriving the Optimum of the Entropy-Regularized Reward Maximization Objective
$=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y|x)}\left[\log\frac{\pi(y|x)}{\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right)}-\log Z(x)\right]$ (20)
@@
which is a valid probability distribution as $\pi^{*}(y|x)\geq 0$ for all $y$ and $\sum_{y}\pi^{*}(y|x)=1$. Since $Z(x)$ is not a function of $y$, we can then re-organize the final objective in Eq D.1 as:
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{E}_{y\sim\pi(y|x)}\left[\log\frac{\pi(y|x)}{\pi^{*}(y|x)}\right]-\log Z(x)\right]=$ (21)
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{D}_{\text{KL}}(\pi(y|x)\mid\mid\pi^{*}(y|x))-\log Z(x)\right]$ (22)
Since $Z(x)$ is independent of $\pi$, the minimum is attained by the policy that minimizes the first KL term. By Gibbs’ inequality, the KL divergence reaches its minimum value of zero if and only if the two distributions are identical. Therefore, this yields the optimal solution. :
@@
D.2 Deriving the SimPO Objective Under the Bradley-Terry Model
@@
Substituting Eq. 25 into Eq. 24 we obtain:
$p^{*}(y_{1}\succ y_{2}|x)$ $=\frac{\exp\left(\alpha\log\pi^{*}(y_{1}|x)+\alpha\log Z(x)\right)}{\exp\left(\alpha\log\pi^{*}(y_{1}|x)+\alpha\log Z(x)\right)+\exp\left(\alpha\log\pi^{*}(y_{2}|x)+\alpha\log Z(x)\right)}$
$=\frac{1}{1+\exp\left(\alpha\log\pi^{*}(y_{2}|x)-\alpha\log\pi^{*}(y_{1}|x)\right)}$
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right).$
@@
D.3 Deriving the SimPO Objective Under the Plackett-Luce Model
The Plackett-Luce model (Plackett, 1975) extends the Bradley-Terry model from pairwise comparisons to full rankings. As in the Bradley-Terry framework, the probability of selecting an option is assumed to be proportional to the value of an underlying latent reward function. In our setting, given a prompt $x$ and a collection of $K$ candidate answers $y_{1},\ldots,y_{K}$, the user produces a permutation $\tau:[K]\to[K]$ that represents their ranking of the answers. Under the Plackett-Luce model, the probability of such a ranking is defined as follows:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp(r^{*}(x,y_{\tau(k)}))}{\sum_{j=k}^{K}\exp(r^{*}(x,y_{\tau(j)}))}$ (26)
Observe that when $K=2$, Equation 26 simplifies to the Bradley-Terry model. For the general Plackett-Luce model, however, we can still leverage the reward parameterization by substituting the reward function expressed in terms of its optimal policy. As in Appendix D.2, the normalization constant $Z(x)$ cancels out, leaving us with:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp\left(\alpha\log\pi^{*}(y_{\tau(k)}|x)\right)}{\sum_{j=k}^{K}\exp\left(\alpha\log\pi^{*}(y_{\tau(j)}|x)\right)}$ (27)
Similarly to the approach for standard DPO, if we have access to a dataset $\mathcal{D}=\{\tau^{(i)},y_{1}^{(i)},\ldots,y_{K}^{(i)},x^{(i)}\}_{i=1}^{N}$ of prompts and user-specified rankings, we can use a parameterized model and optimize this objective with maximum-likelihood.:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{\tau,y_{1},\ldots,y_{K},x\sim\mathcal{D}}\left[\log\prod_{k=1}^{K}\frac{\exp\left(\alpha\log\pi_{\theta}(y_{\tau(k)}|x)\right)}{\sum_{j=k}^{K}\exp\left(\alpha\log\pi_{\theta}(y_{\tau(j)}|x)\right)}\right]$ (28)
D.4 Deriving the Gradient of the SimPO Objective
@@
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\nabla_{\theta}\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\alpha\log\pi_{\theta}(y_{w}|x)-\alpha\log\pi_{\theta}(y_{l}|x)\right)\right]$ (29)
We can rewrite the RHS of Equation 29 as
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\frac{\sigma^{\prime}\left(u\right)}{\sigma\left(u\right)}\nabla_{\theta}\left(u\right)\right],$ (30)
@@
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=\\ -\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg[\alpha\sigma\left(\alpha\log\pi_{\theta}(y_{l}|x)-\alpha\log\pi_{\theta}(y_{w}|x)\right)\bigg[\nabla_{\theta}\log\pi(y_{w}\mid x)-\nabla_{\theta}\log\pi(y_{l}\mid x)\bigg]\bigg],$
@@
D.5 Proof of Lemma 1 and 2 from DPO for Maximum Entropy RL
@@
We say that two reward functions $r(x,y)$ and $r^{\prime}(x,y)$ are from the same equivalence class if $r^{\prime}(x,y)=r(x,y)+f(x)$ for some function $f$. We consider the general Plackett-Luce (with the Bradley-Terry model a special case for $K=2$) and denote the probability distribution over rankings induced by a particular reward function $r(x,y)$ as $p_{r}$. For any prompt $x$, answers $y_{1},\ldots,y_{K}$ and ranking $\tau$ we have:
$p_{r^{\prime}}(\tau|y_{1},\ldots,y_{K},x)$ $=\prod_{k=1}^{K}\frac{\exp(r^{\prime}(x,y_{\tau(k)}))}{\sum_{j=k}^{K}\exp(r^{\prime}(x,y_{\tau(j)}))}$
$=\prod_{k=1}^{K}\frac{\exp(r(x,y_{\tau(k)})+f(x))}{\sum_{j=k}^{K}\exp(r(x,y_{\tau(j)})+f(x))}$
$=\prod_{k=1}^{K}\frac{\exp(f(x))\exp(r(x,y_{\tau(k)}))}{\exp(f(x))\sum_{j=k}^{K}\exp(r(x,y_{\tau(j)}))}$
$=\prod_{k=1}^{K}\frac{\exp(r(x,y_{\tau(k)}))}{\sum_{j=k}^{K}\exp(r(x,y_{\tau(j)}))}$
$=p_{r}(\tau|y_{1},\ldots,y_{K},x),$
@@
Let us consider two reward functions from the same class, such that $r^{\prime}(x,y)=r(x,y)+f(x)$ and, let us denote as $\pi_{r}$ and $\pi_{r^{\prime}}$ the corresponding optimal policies. For all $x,y$ we have
$\pi_{r^{\prime}}(y|x)$ $=\frac{1}{\sum_{y}\exp\left(\frac{1}{\alpha}r^{\prime}(x,y)\right)}\exp\left(\frac{1}{\alpha}r^{\prime}(x,y)\right)$
$=\frac{1}{\sum_{y}\exp\left(\frac{1}{\alpha}(r(x,y)+f(x))\right)}\exp\left(\frac{1}{\alpha}(r(x,y)+f(x))\right)$
$=\frac{1}{\exp\left(\frac{1}{\alpha}f(x)\right)\sum_{y}\exp\left(\frac{1}{\alpha}r(x,y)\right)}\exp\left(\frac{1}{\alpha}r(x,y)\right)\exp\left(\frac{1}{\alpha}f(x)\right)$
$=\frac{1}{\sum_{y}\exp\left(\frac{1}{\alpha}r(x,y)\right)}\exp\left(\frac{1}{\alpha}r(x,y)\right)$
@@
which completes the proof. ∎
D.6 Proof of Theorem 1 from DPO for Maximum Entropy RL
@@
$f(r,\alpha)(x,y)=\alpha\log\pi_{r}(y|x)=\alpha\log\pi_{r^{\prime}}(y|x)=f(r^{\prime},\alpha)(x,y)$
@@
$r^{\prime}(x,y)=r(x,y)+f(x)=\alpha\log\pi(y|x)+f(x)=\alpha\log\pi(y|x)\exp(\frac{1}{\alpha}f(x))=\alpha\log\pi^{\prime}(y|x)$

### v2

@@ Appendix B Margins and Overoptimization
It has been shown that methods such as SimPO can achieve performance comparable to DPO even with a target margin of $\gamma=0$, as demonstrated in the original SimPO paper. This suggests that offline methods do not necessarily require reference models when operating within the safe KL region, and that introducing margins generally improves performance across benchmarks. This effect arises from both model capabilities and dataset coverage: larger models are less prone to common overfitting behaviors and can extract more meaningful signals during optimization, rather than engaging in reward hacking, a phenomenon observed in both online and offline preference optimization (gao2022scalinglawsrewardmodel; rafailov2024directpreferenceoptimizationlanguage). Consequently, the influence of the reference model is minimal and can often be neglected. However, this behavior is contingent on the task being sufficiently challenging and the model being strong enough to avoid overoptimization. To validate this observation, we train Pythia-1B on TL;DR using SimPO across different margin values ($\gamma$) and learning rates, in a setting where the model is relatively weaker and the task is easier compared to standard chat datasets such as UltraFeedback (cui2024ultrafeedbackboostinglanguagemodels) used in SimPO.
@@
Appendix C Mathematical Derivations for Maximum Entropy RL
C.1 Deriving the Optimum of the Entropy-Regularized Reward Maximization Objective
In this appendix, we will derive the optimal policy for Maximum Entropy RL. Analogously to the KL-constrained case (rafailov2024directpreferenceoptimizationlanguage), we optimize the following objective:
@@ C.1 Deriving the Optimum of the Entropy-Regularized Reward Maximization Objective
$=\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{y\sim\pi(y|x)}\Bigg[\log\frac{\pi(y|x)}{\frac{1}{Z(x)}\exp\left(\frac{1}{\alpha}r(x,y)\right)}$
$\qquad\qquad\qquad\qquad-\log Z(x)\Bigg]$
@@
which is a valid probability distribution as $\pi^{*}(y|x)\geq 0$ for all $y$ and $\sum_{y}\pi^{*}(y|x)=1$. Since $Z(x)$ is not a function of $y$, we can then re-organize the final objective in Eq. C.1 as:
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\bigg[\mathbb{E}_{y\sim\pi(y|x)}\left[\log\frac{\pi(y|x)}{\pi^{*}(y|x)}\right]$
$\qquad\qquad\qquad-\log Z(x)\bigg]=$
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[D_{\text{KL}}(\pi(y|x)\|\pi^{*}(y|x))-\log Z(x)\right]$
Since $Z(x)$ is independent of $\pi$, the minimum is attained by the policy that minimizes the first KL term. By Gibbs’ inequality, the KL divergence reaches its minimum value of zero if and only if the two distributions are identical. Therefore, this yields the optimal solution:
@@
C.2 Deriving the SimPO Objective Under the Bradley-Terry Model
@@
Substituting Eq. C.2 into Eq. C.2 we obtain:
$p^{*}(y_{1}$ $\succ y_{2}|x)$
$=\frac{\exp(\alpha\log\pi^{*}(y_{1}|x))}{\sum_{i\in\{1,2\}}\exp(\alpha\log\pi^{*}(y_{i}|x))}$
$=\frac{1}{1+\exp\left(\alpha\log\frac{\pi^{*}(y_{2}|x)}{\pi^{*}(y_{1}|x)}\right)}$
$=\sigma\left(\alpha\log\pi^{*}(y_{1}|x)-\alpha\log\pi^{*}(y_{2}|x)\right)$
@@
C.3 Deriving the SimPO Objective Under the Plackett-Luce Model
The Plackett-Luce model (af5079a1-8ca5-3727-a405-0a82390327b7) extends the Bradley-Terry model from pairwise comparisons to full rankings. As in the Bradley-Terry framework, the probability of selecting an option is assumed to be proportional to the value of an underlying latent reward function. In our setting, given a prompt $x$ and a collection of $K$ candidate answers $y_{1},\ldots,y_{K}$, the user produces a permutation $\tau:[K]\to[K]$ that represents their ranking of the answers. Under the Plackett-Luce model, the probability of such a ranking is defined as follows. Let $r_{k}=r^{*}(x,y_{\tau(k)})$. Then:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp(r_{k})}{\sum_{j=k}^{K}\exp(r_{j})}$ (11)
Observe that when $K=2$, Equation 11 simplifies to the Bradley-Terry model. For the general Plackett-Luce model, however, we can still leverage the reward parameterization by substituting the reward function expressed in terms of its optimal policy. As in Appendix C.2, the normalization constant $Z(x)$ cancels out. Let $s_{k}=\alpha\log\pi^{*}(y_{\tau(k)}|x)$. Then:
$p^{*}(\tau|y_{1},\ldots,y_{K},x)=\prod_{k=1}^{K}\frac{\exp(s_{k})}{\sum_{j=k}^{K}\exp(s_{j})}$ (12)
Similarly to the approach for standard DPO, if we have access to a dataset $\mathcal{D}=\{\tau^{(i)},y_{1}^{(i)},\ldots,y_{K}^{(i)},x^{(i)}\}_{i=1}^{N}$ of prompts and user-specified rankings, we can use a parameterized model and optimize this objective with maximum-likelihood. Let $s_{k}^{\theta}=\alpha\log\pi_{\theta}(y_{\tau(k)}|x)$. Then:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})$
$=-\mathbb{E}_{\mathcal{D}}\left[\sum_{k=1}^{K}\log\frac{\exp(s_{k}^{\theta})}{\sum_{j=k}^{K}\exp(s_{j}^{\theta})}\right]$
C.4 Deriving the Gradient of the SimPO Objective
@@
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})$
$=-\nabla_{\theta}\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\alpha\log\pi_{\theta}(y_{w}|x)-\alpha\log\pi_{\theta}(y_{l}|x)\right)\right]$
We can rewrite the RHS of Equation C.4 as
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\frac{\sigma^{\prime}\left(u\right)}{\sigma\left(u\right)}\nabla_{\theta}\left(u\right)\right],$
@@
$\nabla_{\theta}\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\\ \bigg[\alpha\sigma\left(\alpha\log\pi_{\theta}(y_{l}|x)-\alpha\log\pi_{\theta}(y_{w}|x)\right)\\ \bigg[\nabla_{\theta}\log\pi(y_{w}\mid x)-\nabla_{\theta}\log\pi(y_{l}\mid x)\bigg]\bigg],$
@@
C.5 Proof of Lemma 1 and 2 from DPO for Maximum Entropy RL
@@
We say that two reward functions $r(x,y)$ and $r^{\prime}(x,y)$ are from the same equivalence class if $r^{\prime}(x,y)=r(x,y)+f(x)$ for some function $f$. We consider the general Plackett-Luce (with the Bradley-Terry model a special case for $K=2$) and denote the probability distribution over rankings induced by a particular reward function $r(x,y)$ as $p_{r}$. For any prompt $x$, answers $y_{1},\ldots,y_{K}$ and ranking $\tau$ we have: Let $r_{k}=r(x,y_{\tau(k)})$ and $r^{\prime}_{k}=r^{\prime}(x,y_{\tau(k)})=r_{k}+f(x)$. Then: Let $r_{k}=r(x,y_{\tau(k)})$ and $r^{\prime}_{k}=r_{k}+f(x)$. Then:
$p_{r^{\prime}}$ $(\tau|y_{1},\ldots,y_{K},x)$
$=\prod_{k=1}^{K}\frac{\exp(r^{\prime}_{k})}{\sum_{j=k}^{K}\exp(r^{\prime}_{j})}$
$=\prod_{k=1}^{K}\frac{\exp(r_{k}+f(x))}{\sum_{j=k}^{K}\exp(r_{j}+f(x))}$
$=\prod_{k=1}^{K}\frac{\exp(f(x))\exp(r_{k})}{\exp(f(x))\sum_{j=k}^{K}\exp(r_{j})}$
$=\prod_{k=1}^{K}\frac{\exp(r_{k})}{\sum_{j=k}^{K}\exp(r_{j})}$
$=p_{r}(\tau|y_{1},\ldots,y_{K},x)$
@@
Let us consider two reward functions from the same class, such that $r^{\prime}(x,y)=r(x,y)+f(x)$ and, let us denote as $\pi_{r}$ and $\pi_{r^{\prime}}$ the corresponding optimal policies. For all $x,y$, let $\rho=\frac{1}{\alpha}r(x,y)$, $\rho^{\prime}=\frac{1}{\alpha}r^{\prime}(x,y)$, and $\phi=\frac{1}{\alpha}f(x)$. Then:
$\pi_{r^{\prime}}(y|x)$ $=\frac{\exp(\rho^{\prime})}{\sum_{y}\exp(\rho^{\prime})}$
$=\frac{\exp(\rho+\phi)}{\sum_{y}\exp(\rho+\phi)}$
$=\frac{\exp(\phi)\exp(\rho)}{\exp(\phi)\sum_{y}\exp(\rho)}$
$=\frac{\exp(\rho)}{\sum_{y}\exp(\rho)}$
@@
C.6 Proof of Theorem 1 from DPO for Maximum Entropy RL
@@
$f(r,\alpha)(x,y)=\alpha\log\pi_{r}(y|x)=$
$\alpha\log\pi_{r^{\prime}}(y|x)=f(r^{\prime},\alpha)(x,y)$
@@
$r^{\prime}(x,y)=r(x,y)+f(x)=\alpha\log\pi(y|x)+f(x)$
$=\alpha\log\pi(y|x)\exp(\frac{1}{\alpha}f(x))$
$=\alpha\log\pi^{\prime}(y|x)$
@@
Appendix D Extra Figures
Figure 10: Reward dynamics and KL divergence metrics for entropy-regularized RL training across different entropy coefficients. Top-left panel shows reward progression (RLHF reward) over training steps for various entropy values. Top-right panel shows KL divergence between the current policy and the SFT reference policy ($KL(\pi_{t}||\pi_{\text{SFT}})$). Bottom-left panel tracks entropy reward across training steps. Bottom-right panel displays KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All plots are based on the Pythia-6.9B model trained with the learning rate of $1\times 10^{-6}$.
Figure 11: Reward dynamics and KL divergence metrics for entropy-regularized RL training across different entropy coefficients. Top-left panel shows reward progression (RLHF reward) over training steps for various entropy values. Top-right panel shows KL divergence between the current policy and the SFT reference policy ($KL(\pi_{t}||\pi_{\text{SFT}})$). Bottom-left panel tracks entropy reward across training steps. Bottom-right panel displays KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All plots are based on the Pythia-6.9B model trained with the learning rate of $1\times 10^{-7}$.
Figure 12: Training metrics for KL-constrained RL on the Pythia-6.9B model. Top panel shows the KL divergence between the policy and reference SFT policy ($KL(\pi_{t}||\pi_{\text{SFT}})$) over training steps. Middle panel displays the reward trajectory (RLHF reward). Bottom panel shows the KL divergence between consecutive policy updates ($KL(\pi_{t}||\pi_{t-1})$). All results correspond to a single training run with a fixed KL constraint.